OpenStudy (anonymous):

Derivative Help! f(x)=sqroot(1-2x) I will fill in what I have done

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so you know f(x) = (1-2x)^(1/2) (apply the power rule to get (1/2)(1-2x)^(-1/2)(-2) make it cleaner to obtain -(1-2x)^(-1/2) (get rid of the negative exponents to finally obtain f'(x) = -1/(sqrt(1-2x) :)

OpenStudy (anonymous):

compare it to your answer

OpenStudy (anonymous):

\[f \prime= \lim_{h \rightarrow 0} \frac{ \sqrt{1-2(x+h)}-\sqrt{1-2x} }{ h } * \frac{ \sqrt{1-2(x+h)}+\sqrt{1-2x} }{ \sqrt{1-2(x+h)}+\sqrt{1-2x} }=\frac{ -2 }{ \sqrt{1-2(x+h)}+\sqrt{1-2x}}\]

OpenStudy (anonymous):

it cut it off hang on I am stuck at this point

OpenStudy (anonymous):

your good man

OpenStudy (anonymous):

but

OpenStudy (anonymous):

this is where i get stuck \[\lim_{h \rightarrow 0}\frac{ -2 }{ \sqrt{1-2(x+h)}+\sqrt{1-2x} }\]

OpenStudy (anonymous):

yeah your perfect sub h as 0 now

OpenStudy (anonymous):

because this is a limit problem ans since the last 2 entries are the same 2 at the bottom and the top will cancel out and you have my answer :)

OpenStudy (anonymous):

got it i guess i was trying to make it harder than it needed to be thanks